A 2500 kg car accelerates from rest to a velocity of 30.0 m/s in 8.00 seconds. The work done by the engine to accelerate the car is Multiple Choice 3.10 × 106 J. 2.56 × 106 J. 2.01 × 106 J. 1.85 × 106 J. 1.13 × 106 J.  
Answer #1

____________________________________________________

Acceleration = a = (30-0)/8 = 3.75 m/s2
Now,
v2 = u2 + 2as
302 = 0 + 2 x 3.75 x s
Distance = s= 120 meters
Work = F.s = 2500 x 3.75 x 120 = 1.13 x 106 Joules
___________________________________________________

Please rate
if any mistake in this answer please comment i will clarify your doubt . thank you